The number of ordered pair (a,b), where a,b E (-pi,pi),


satisfying cos(a-b) = 1 and cos(a + b) = 1/e ,is.......


(E means 'belongs to')


 


pls explain.....

Grade:12

2 Answers

Amitabh Shrivastava
12 years ago

cos(a+b)=1/e => a+b=+or-cosinverse(1/e)

cos(a-b)=1=>a=bORa-b=pi  the latter does not satisfy cos(a+b)=1/e

hence a=b;

Thus a=b=+OR-cosinverse(1/e)/2

Thus two ordered pairs

Parnav Bansal
5 years ago
Plz give this question answer properly …
Above answer is not correct explaination.....,.........................